OpenStudy (anonymous):

Parallel, perpendicular, or neither: y=4x-3 y=2x-3

OpenStudy (baldymcgee6):

What do the slopes and parallel/perpendicular have to do with each other? Is there a way we can recognize this?

OpenStudy (anonymous):

?

OpenStudy (baldymcgee6):

If you have two lines that are parallel, they have the same slope... right?

OpenStudy (anonymous):

If you graph the two you can see that they intersect at point (0,-3) so neither.

OpenStudy (anonymous):

correct but with these two equations the slope is double on the first one > 4x-3

OpenStudy (baldymcgee6):

@sjerman1, i'm trying to make a point here. OBVIOUSLY the two functions have different slopes, therefore they are not parallel. Btw there is no need to graph the two functions. Perpendicular slopes can be recognized when you see the slope is the negative reciprocal of the other slope. Eg, 2 and -1/2 As we can see just by looking at the slope of the two functions, (not the graph), they are neither perpendicular or parallel.

OpenStudy (anonymous):

ok thank you so. How can you recognize if they are perpendicular.

OpenStudy (baldymcgee6):

Perpendicular slopes can be recognized when you see the slope is the negative reciprocal of the other slope. Eg, 2 and -1/2

OpenStudy (anonymous):

so y=2x+4 y=-1/2x+4 would be perpendicular then right

OpenStudy (baldymcgee6):

Perfect.

OpenStudy (baldymcgee6):

Just remember: Parallel = same slope Perpendicular = negative reciprocal slope

OpenStudy (anonymous):

So we are only looking at the slope and not the + or - 4

OpenStudy (baldymcgee6):

Linear function (lines) are in the form of y = mx+b, where m is the slope and b is the y intercept. y=2x+4 Slope: 2 y intercept: 4 y=-1/2x+4 Slope: -1/2 y int: 4

OpenStudy (anonymous):

Yes, so for perpendicular lines do we only look at the slope or do we have to look at the y intercept also?

OpenStudy (anonymous):

Only consider the number and sign in front of x when equation is written as y = *x +/- b